What is the solubility of zinc hydroxide in a buffer with pH=10


KSp is 5.0E-17
Answer #1

use:

pH = -log [H+]

10 = -log [H+]

[H+] = 1*10^-10 M

use:

[OH-] = Kw/[H+]

Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC

[OH-] = (1.0*10^-14)/[H+]

[OH-] = (1.0*10^-14)/(1*10^-10)

[OH-] = 1*10^-4 M

At equilibrium:

Zn(OH)2 <----> Zn2+ + 2 OH-

   s 1*10^-4 + 2s

Ksp = [Zn2+][OH-]^2

5*10^-17=(s)*(1*10^-4+ 2s)^2

Since Ksp is small, s can be ignored as compared to 1*10^-4

Above expression thus becomes:

5*10^-17=(s)*(1*10^-4)^2

5*10^-17= (s) * 10^-8

s = 5.0*10^-9 M

Answer: 5.0*10^-9 M